OpenStudy (calculator):

Solve 2cos(x)-4sin(x)=3 [0,360]

OpenStudy (anonymous):

cos(x)^2 + sin(x)^2 = 1 so cos(x) = [1 -sin(x)^2]^0.5 plug it into the equation: 2[1-sin(x)^2]^0.5 - 4sin(x) = 3 2[1-sin(x)^2]^0.5 = 3 + 4sin(x) square both sides 4[1-sin(x)^2] = 9 + 24sin(x) + 16sin(x)^2 can you do it now ?

OpenStudy (calculator):

can you do it using harmonic form

OpenStudy (cwrw238):

another way to do this is by complementary angle method angle A R cos (x + A) = R cosx cosA - R sinx x sin A = 1

OpenStudy (anonymous):

mm Rcos(x+a) = Rcos(x)cos(a) - Rsin(x)sin(a) Rcos(a) = 2 , Rsin(a) = 4 -> tan(a) = 2 -> a = 63.43 so R = 4.47

OpenStudy (calculator):

I found \[2\sqrt{5}\cos(\theta+116.6)-3=0\]

OpenStudy (calculator):

\[\text{this is how i find} \\ \\ 2\cos \theta-4\sin \theta=3~~~~~~~~~Rcos(\theta+\alpha) \\ \\ 2\cos \theta-4\sin \theta -3=0 ~~~~~~~Rcos \theta \cos \alpha - R \sin \theta \sin \alpha \\ \\ \cos \theta (2-Rcos \alpha)-\sin \theta (4+\sin \alpha ) -3=0 \\ \\ Rcos \alpha=2 ~~~~~~~Rsin \alpha=-4 \\ \\ \tan \alpha =-2 \\ \\ \alpha=116.6 \\ \\R=\sqrt{4+16} \\ \\R=2\sqrt{5} \\ \\ 2\sqrt{5}\cos( \theta +116.6)-3=0\]

OpenStudy (amorfide):

oh crap i made a mistake... i carried on oops 2cos(x)-4sin(x)=3 2cos(x)-4sin(x) is equivalent to Rcos(x+a) tana=4/2=2 a=tan^-1(2) R=root(a²+b²) R=root 20 R=2root5 2root5Cos(x+63.4)=3 now divide by 2root5 on both sides cos(x+63.4)=3/2root5 inverse cos x+63.4=cos^-1(3/2root5) that should help you

OpenStudy (calculator):

are you sure the angle is 63.4

OpenStudy (amorfide):

Tan(a)=-b/a b=-4 a=2 --4/2=4/2=2 tanA=2 inverse tan(2)=63.4

OpenStudy (amorfide):

does that help you?